OpenStudy (anonymous):

find integral of 1/x(logx)^m dx,m>o does this question have 2 answers?

OpenStudy (lgbasallote):

why 2 answers?

OpenStudy (anonymous):

m=1 and m=2toinfinity

OpenStudy (lgbasallote):

let \[u = \log x\] \[du = \frac 1x dx\] does that help?

OpenStudy (lgbasallote):

and no that's not a limit... m>0 is just to denote m exists

OpenStudy (lgbasallote):

your answer should be in algebra form

OpenStudy (anonymous):

ya i know that but when m=1 its logu,otherwise u need to use the other law...x^n...

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

oh well then yes you ahve 2 answers

OpenStudy (anonymous):

Do this............ find the integral of 1/x^3-x^4

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

\[\int \frac{1}{x^3 } - x^4\] ???

OpenStudy (anonymous):

no x4 is in the denominator along wid x^3

OpenStudy (anonymous):

the one which u thought is damn easy

OpenStudy (anonymous):

x^4

OpenStudy (lgbasallote):

\[\huge \int \frac 1{x^2 - x^4} \implies \int \frac 1{x^3(1 - x)}\] does that help?

OpenStudy (anonymous):

yes then...

OpenStudy (lgbasallote):

uhh that should be 1/(x^3 - x^4) you know what i mean...

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

sry,..

OpenStudy (lgbasallote):

now you use partil fractions... \[\Large \implies \frac Ax + \frac B{x^2} + \frac C{x^3} + \frac D{1-x} = \frac{1}{x^3(1-x)}\] right?

OpenStudy (anonymous):

i did it some other way...lol maybe thats also correct ..that is defintely correct

OpenStudy (lgbasallote):

what did you do?

OpenStudy (lgbasallote):

integration by parts?

OpenStudy (anonymous):

i put the N=1=(1-x)+x,then split ..then PFM

OpenStudy (anonymous):

partial fraction(PFM)

OpenStudy (lgbasallote):

hmm i guess if it works then it's the same

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

im not familiar with that kind of manipulation

OpenStudy (anonymous):

now this one...find integral of x*whole root of 1-x/1+x

OpenStudy (lgbasallote):

\[\int x \sqrt{\frac{1-x}{1+x}}\] ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how do u type like that?

OpenStudy (lgbasallote):

i have no idea

OpenStudy (lgbasallote):

the integral i mean

OpenStudy (anonymous):

k i tried then...i multiplied N and D by the N...

OpenStudy (lgbasallote):

as for my typing...do you see that Equation button below the reply box?

OpenStudy (anonymous):

nah!

OpenStudy (anonymous):

ok yes i can see..

OpenStudy (lgbasallote):

i think what you did is right

OpenStudy (anonymous):

r u a professor or anythng like that?

OpenStudy (lgbasallote):

yes. something like that

OpenStudy (lgbasallote):

anyway...if you multiply N and D by N you get \[\int \frac{x(1-x)}{\sqrt{1-x^2}}\] right?

OpenStudy (anonymous):

but then what? shud i put N=A(d/dxD)+B?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

split it up now \[\int \frac{x}{\sqrt{1-x^2}} + \int \frac{x^2}{\sqrt{1-x^2}}\] first one is solveable by u-sub second one is solveable by trig sub

OpenStudy (lgbasallote):

right?

OpenStudy (anonymous):

oh ya!!!!!!111111!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

btw did u solve that question.............-->find integral of 1+cosxcosA/cosA+cosx ?

OpenStudy (anonymous):

where A is constant

OpenStudy (lgbasallote):

\[\int \frac{1+\cos x \cos A}{\cos A + \cos x}\] ???

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

...well you can try splittig it up again...

OpenStudy (anonymous):

what i did was i put 1=sin^2x+cos^2x

OpenStudy (anonymous):

took cosx common and split

OpenStudy (lgbasallote):

hmm clever

OpenStudy (anonymous):

but then it becomes s^2x/coA+cosx

OpenStudy (anonymous):

s means sine

OpenStudy (lgbasallote):

this is a hard one

OpenStudy (anonymous):

ya....these were the ones i didnt understand from the 400 sums

OpenStudy (lgbasallote):

sorry. i have no idea what this is

OpenStudy (anonymous):

np

OpenStudy (lgbasallote):

good questions though

OpenStudy (anonymous):

ya ........................HOW to do this question here integrate (3sinx-4)cosx/(5sinx+3)^8 w.r.t.x

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

there is dx after w.r.t.x

OpenStudy (lgbasallote):

hmm ho many of these do you have

OpenStudy (anonymous):

before*

OpenStudy (anonymous):

this might be the lst one

OpenStudy (anonymous):

last*

OpenStudy (lgbasallote):

let u = sinx du = cosx dx \[\implies \int \frac{(3u-4)du}{(5u+3)^8}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now by PFM?

OpenStudy (anonymous):

:O

OpenStudy (lgbasallote):

no... i think you should just split the fraction

OpenStudy (lgbasallote):

\[\implies \int\frac{3u}{(5u+3)^8} - \int \frac{4}{(5u+3)^8}\] maybe?

OpenStudy (lgbasallote):

ugh my brain hurts now o.O

OpenStudy (anonymous):

lol